Given the quadratic function f(x) = 25x^2 – 100x + 100, find a value of x such that f(x) = 25.
Answer by Alan3354(69443)

Given the quadratic function f(x) = 25x^2 – 100x + 100, find a value of x such that f(x) = 25.
-----------
25x^2 – 100x + 100 = 25
25x^2 – 100x + 75 = 0
x^2 - 4x + 3 = 0
(x-3)*(x-1) = 0
x = 1, 3